OpenStudy (anonymous):

10x=-80 solve using the multiplication principle

OpenStudy (anonymous):

The multiplication principle says you can multiply both sides of an equation by any nonzero real number, and it will still be true. Our objective is to isolate the variable x. What number do we need to multiply by to have the x standing all by itself?

OpenStudy (anonymous):

10 is that correct?

OpenStudy (anonymous):

Not quite. If we multiply by ten, we get 100x=-800. Do you have another choice?

OpenStudy (anonymous):

i would think 2 because they are both even numbers

OpenStudy (anonymous):

That will give 20x=-160. The number we need to multiply by is the reciprocal of the ten that multiplies the x. Do you know what that is?

OpenStudy (anonymous):

or is it 0?

OpenStudy (anonymous):

Not zero. Zero just wipes out all our information. We need to multiply by 1/10, like this\[10x=-80\implies \frac{1}{10}10x=-80\frac{1}{10}\implies x=-8\]

OpenStudy (anonymous):

Multiplying by the reciprocal is the same as dividing by the number. That is probably the way you are thinking about it.

OpenStudy (anonymous):

10x=-80 x=-80/10 x=-8